plEase help!urgent!
1. An alloy contains 20% gold and another contains 40% gold. How much of each must be mixed to produce 200kg of 25%gold?

2. A dairy man wishes to prepare cream containing 5% butterfat. How much cream of 25% butterfat must be added to 50 kg of cream approximately containing 4% butterfat?

3. Mart is twice as old as Nald. In six years, he will be four times as old as Nald was three years ago. What are their present ages?

4. A freight train traveling at a rate of 60kph leaves a city. Three hours later, an express train whose average rate is 80kph leaves the same city traveling in the same direction. How long will it take the express train to overtake the freight train?

Answer by Alan3354(69443)

1. An alloy contains 20% gold and another contains 40% gold. How much of each must be mixed to produce 200kg of 25%gold?
200 kg @ 25% = 50 kg of gold
x = amount of 20% alloy
x*0.2 + (200-x)*0.4 = 50
2x + 4(200-x) = 500
2x + 800 - 4x = 500
-2x = -300
x = 150 kg of 20% alloy
50 kg of 40% alloy
----------------------
2. A dairy man wishes to prepare cream containing 5% butterfat. How much cream of 25% butterfat must be added to 50 kg of cream approximately containing 4% butterfat?
50*0.04 + x*0.25 = (50+x)*0.05
50*4 + 25x = (50+x)*5
Solve for x.
------------------
3. Mart is twice as old as Nald. In six years, he will be four times as old as Nald was three years ago. What are their present ages?
m = 2n
m+6 = 4(n-3)
2n+6 = 4n-12
-2n = -18
Nald is 9
Mart is 18
------------
4. A freight train traveling at a rate of 60kph leaves a city. Three hours later, an express train whose average rate is 80kph leaves the same city traveling in the same direction. How long will it take the express train to overtake the freight train?
---
In 3 hours the freight train goes 180 km.
The fast train gains on it at 80-60, or 20 kph
180/20 = 9 hours to overtake it.